OpenStudy (anonymous):

Calculate ∂f/∂s+ ∂f/∂t at s = 2, t = −1, given that f=f(x,y); x=s−t, y=s2+t2, ∂f/∂x(3,5)=0.06170, ∂f/∂y(3,5)=0.06170.

OpenStudy (anonymous):

∂f/∂s= ∂f/∂x dx/ds +∂f/∂y dy/ds

OpenStudy (anonymous):

∂f/∂s= ∂f/∂x dx/ds +∂f/∂y dy/ds 0.06170 1 +0.06170 2s

OpenStudy (anonymous):

∂f/∂t= ∂f/∂x dx/dt +∂f/∂y dy/dt 0.06170 ( -1) +.0617 2t

OpenStudy (zarkon):

just notation, but we should write \[\frac{\partial f}{\partial s}=\frac{\partial f}{\partial x}\frac{\partial x}{\partial s}+\frac{\partial f}{\partial y}\frac{\partial y}{\partial s}\] since x and y are functions of more than one variable

OpenStudy (turingtest):

now do you get the partial symbols?

OpenStudy (zarkon):

\partial

OpenStudy (turingtest):

\[/\partial\]thanks!

OpenStudy (anonymous):

ok thanks for the help... just one more question. how do you know that ∂f/∂s= ∂f/∂x?

OpenStudy (anonymous):

nvm, i got it now. thanks